# HDC: Hierarchical Semantic Decoding with Counting Assistance for Generalized Referring Expression Segmentation

### v1

@@ Abstract
The newly proposed Generalized Referring Expression Segmentation (GRES) amplifies the formulation of classic RES by involving multiple/non-target scenarios. Recent approaches focus on optimizing the last modality-fused feature which is directly utilized for segmentation and object-existence identification. However, the attempt to integrate all-grained information into a single joint representation is impractical in GRES due to the increased complexity of the spatial relationships among instances and deceptive text descriptions. Furthermore, the subsequent binary target justification across all referent scenarios fails to specify their inherent differences, leading to ambiguity in object understanding. To address the weakness, we propose a Hierarchical Semantic Decoding with Counting Assistance framework (HDC). It hierarchically transfers complementary modality information across granularities, and then aggregates each well-aligned semantic correspondence for multi-level decoding. Moreover, with complete semantic context modeling, we endow HDC with explicit counting capability to facilitate comprehensive object perception in multiple/single/non-target settings. Experimental results on gRefCOCO, Ref-ZOM, R-RefCOCO, and RefCOCO benchmarks demonstrate the effectiveness and rationality of HDC which outperforms the state-of-the-art GRES methods by a remarkable margin. Code will be available here.
@@ 1 Introduction
However, it poses new challenges with the additional generalized settings. The multiple-target referent necessitates more precise semantic clues to handle complicated spatial-text relationships between instances for arbitrary location segmentation. On the other hand, the empty-target one highlights the significance of contextual information in mitigating undesirable responses from similar objects. Therefore, an intuitive approach is to endow prevalent RES paradigms [33, 6, 7] with new capability. For instance, ReLA [20] introduces a large-scale generalized dataset gRefCOCO and utilizes spatial-aware queries to transform GRES into a related region retrieval task. Meanwhile, DMMI [11] and RefSegformer [30] enhance modality interaction in the early-fusion process and incorporate new benchmarks Ref-ZOM and R-RefCOCO respectively, to realize robust segmentation.
Figure 1: Illustration of different decoding processes. Previous GRES methods attempt to integrate all-grained context into one joint feature which is directly for mask and object-existence prediction (a), exhibiting vulnerability to challenges of multiple/non-target scenarios (c). Instead of relying on the single representation, we perform hierarchical semantic decoding with multi-level visual-linguistic correspondence aggregation for robust segmentation. Moreover, in contrast to simple binary classification, we empower HDC with counting ability to promote overall object awareness (b).
Although these methods have made advancements in GRES, they still have intrinsic weaknesses in two aspects: 1) They are limited by the bottleneck of capturing complete associations between visual and linguistic contents. As Fig. 1 shows, they strive to optimize the last modality-fused feature which is directly for final mask prediction. However, the extra complexity of text descriptions and geometric relationships among instances in GRES impedes the integration of all-grained contexts into one joint representation. Without complete modality alignment at each granularity, it potentially leads to incorrect or incomplete object response. 2) A simple binary classification head for object-existence identification not only occurs vulnerability to deceptive instances with accumulated bias in single representation but also incurs ambiguity for ignoring inherent differences of one-one/many referents.
Therefore, the significant problem lies in how to excavate the potential of multi-grained semantic correspondence between visual and linguistic when decoding, as well as fully facilitate the comprehensive object perception by considering the specificity of each scenario. Motivated by that, we present a Hierarchical Semantic Decoding with Counting Assistance framework (HDC). In HDC, a Hierarchical Semantic Decoding module (HSD) progressively transfers complementary information between modalities across different granularities to formulate intricate instance-linguist correlation hierarchies. Instead of leveraging single joint features, HSD utilizes multi-level semantic aggregation for robust segmentation. With hierarchical semantic-context alignment, we further introduce Multi-Adaptive Object Counting module (MAC) on HSD to empower the model with explicit counting ability, which improves overall object awareness under each referring setting.
We conduct extensive experiments on three popular GRES benchmarks, i.e., gRefCOCO [20], Ref-ZOM [11] and R-RefCOCO [30] and RES datasets RefCOCO/+/g [15, 27] to validate the effectiveness of our methods. Results show that HDC sufficiently meets the challenges of GRES and addresses the limitations of existing methods, which notably outperforms them for all benchmarks.
@@
  • We present a Hierarchical Semantic Decoding with Counting Assistance framework (HDC) for GRES to boost the multi-grained comprehension of complex referring semantic associations at different levels. In HDC, a Hierarchical Semantic Decoding module (HSD) is designed to progressively transfer mutual modality complementary information across granularities, and utilize each well-aligned semantic correspondence for multi-level decoding.
  • Built on HSD, we further introduce Multi-Adaptive Object Counting module (MAC) to facilitate comprehensive object perception by assigning explicit counting capability, which seamlessly adapts to the specificity of multiple/single/non-target scenarios. To our best knowledge, we are the first to bring object counting into GRES.
  • Without bells and whistles, our approach outperforms all existing state-of-the-art methods by a remarkable margin under fair comparison, e.g., $+4.7\%$ gIoU in gRefCOCO, $+6.7\%$ rIoU in R-RefCOCO and $+5.3\%$ Acc. in Ref-ZOM.
@@ 2 Related Work
Referring Expression Segmentation (RES) aims to localize a single object referred by a natural language description, which is first introduced by Hu et al. [10]. Conventional approaches [10, 22, 1, 12, 7] typically utilize convolution-based operations as the cross-modal fusion to generate segmentation masks. To overcome the weakness of insufficiency in modeling visual-text relation that prior works suffer, recent studies [33, 6, 21, 24, 16] design advanced attention-based mechanism [29] to facilitate multi-modal interaction.
Since previous works are restricted by pre-defined rules that one sentence must match an object, ReLA [20] proposes the Generalized Referring Expression Segmentation (GRES) task, which additionally involves both empty-target and multiple-target scenarios. Moreover, DMMI [11] introduces a new benchmark and baseline to realize beyond one target segmentation. Similarly, RefSegformer [30] endows the transformer-based model with empty-target sentence discrimination, achieving robust segmentation. However, they seek to integrate all-grained contents into a single joint representation for mask prediction, which is impractical in GRES. In contrast, we aggregate visual-linguistic semantic clues in hierarchical structures for more generalized segmentation.
@@
Object counting is to predict the number of conditional objects within an image, which can be divided into detection-based and regression-based approaches. The detection-based methods [13, 8, 9] obtain the number of instances by proposals localization and ranking while the regression-based ones [3, 18, 28] generates predicted counts via density map integration. Different from the classic object counting mechanism, our proposed Multi-Adaptive Object Counting module performs at the query level. By considering various scenarios, i.e., multiple/single/non-target referents, our model achieves comprehensive object perception under each situation.
@@ 3.1 Overview
Figure 2: Overview of HDC. The model takes an image with the referring expression as input. After the encoding process, the Hierarchical Semantic Decoding module (HSD) transfers mutual modality complementary information hierarchically for multi-grained alignment and then aggregates each to perform multi-level decoding for final segmentation. Moreover, with well-aligned semantic contexts, we introduce Multi-Adaptive Object Counting module (MAC) to fully promote comprehensive object perception by the category-specific counting ability in multiple/single/non-target scenarios.
@@
Fig. 2 depicts the overall pipeline of our HDC. Given an image $\mathcal{I}$, the visual backbone, e.g., Swin-Base [25] is to extract a set of vision features $\mathcal{V}_{i}\in\mathrm{R}^{D^{V}\times H_{i}\times W_{i}}$, $i\in\{1,2,3,4\}$, where $H_{i}$ and $W_{i}$ denote the width and height of each scale of feature maps respectively. $D^{V}$ signifies the channel dimension. $\mathcal{V}_{i}$ is further encoded as mask features $\mathcal{V}^{m}_{i}$ by deformable attention [35]. Simultaneously, the textual expression $\mathcal{E}$ with $T$ words is sent into a transformer-based [5] language encoder which generates the word-level embedding $\mathcal{L}^{w}\in\mathbb{R}^{T\times D^{L}}$ and sentence-level embedding $\mathcal{L}^{s}\in\mathbb{R}^{D^{L}}$. Considering that $\mathcal{L}^{w}$ contains fine-grained descriptions while $\mathcal{L}^{s}$ expresses the general information, language query $\mathcal{L}^{O}\in\mathbb{R}^{N\times D^{L}}$ is obtained via the concatenation of them. Meantime, $\mathcal{L}^{O}$ and $\mathcal{V}^{m}_{i}$ are sent into the Hierarchical Semantic Decoding (HSD) module, which generates corresponding object segmentation $\mathcal{M}\in\mathbb{R}^{H\times W}$ by the progressively mutual modality complementary across multi-granularity and multi-level semantic aggregation. In addition, Multi-Adaptive Object Counting (MAC) is designed to promote the precise object perception of different scenarios by category-specific counting ability. We elaborate on each module in the following sections.
3.2 Hierarchical Semantic Decoding Module
The extended formulation of multiple-target and empty-target referents in GRES brings significant challenges since the former necessities a more precise cross-modal understanding of the complex association between objects, e.g., different domains, geometrical relationships while the latter highlights long-range contextual information for mitigating interference from similar deceptive objects. Previous GRES methods [20, 11, 30] struggle to optimize the last modality-fused features for the subsequent mask prediction, ignoring the significance of explicit modality alignment at each granularity. While it is acceptable in classic RES, the extremely intricate referring semantics between instances and text in GRES hinder the integration of all-grained modality correspondence in one representation, thus incurring misinterpretation. To address the issues, we propose a Hierarchical Semantic Decoding module to perform mutual modality calibration at each granularity and multi-level semantic integration for robust segmentation. Specifically, we utilize cascaded Semantic To Mask/Query modules to progressively transfer complementary information between modalities, which captures the high-level visual-linguistic correlation at each hierarchy. Then the subsequent Multi-level Mask Aggregation is to perform semantic aggregation for object-aware segmentation.
Semantic To Mask Module.
@@ 3.2 Hierarchical Semantic Decoding Module
The core idea of STM is to perform bi-directional modality complementary to generate the fine-grained semantic map at each scale of visual features, which is further utilized in the multi-level decoding process. As Fig. 3 shows, the visual feature $\mathcal{V}^{m}_{i}$ and the language query $\mathcal{L}^{O}$ are concurrently input in STM. To be specific, taking $\mathcal{V}^{m}_{1}$ as an example, we firstly leverage fully connected layers to map visual features and language query into a shared joint space $D$, respectively. The mapped representations are further denoted as $\mathcal{V}$ and $\mathcal{L}$ for clarity. Then, we derive the coarse semantic map $\mathcal{A}$ by attention mechanism and it is applied in visual-linguistic interaction to align information flow between modalities. The preliminary modality complementary highlights the corresponding object activation of language-guided visual features, while simultaneously relieving the potential noise in visual-oriented textual representations. The above process can be formulated as:
@@
where $\mathcal{W}^{k}_{V}$, $\mathcal{W}^{v}_{V}$, $\mathcal{W}^{k}_{L}$, $\mathcal{W}^{v}_{L}$ are learnable weights that map visual and linguistic representations into attention space.
As shown in Fig. 3, after generating the object response in $\mathcal{F}^{V\sim L}$ (textual activation of $H_{1}W_{1}$ pixels) and enhancing the informative tokens in $\mathcal{F}^{L\sim V}$ (visual response for $N$ tokens), we perform high-level cross-modal supplement between them to generate the fine-grained semantic map $\mathcal{S}$:
$\mathcal{S}=\mathcal{F}^{L\sim V}\times(\mathcal{F}^{V\sim L})^{T}.$ (4)
In this manner, the key tokens related to the referent are anticipated to emerge significant responses in the semantic map while suppressing the activation of non-related regions. Considering that spatial relationships between instances and empty-target identification are crucial in GRES, solely relying on the last plain visual feature [20, 11] for mask generation potentially causes the misinterpretation of the referred target. Therefore, we provide an effective way to address the issue, which is the direct utilization of the fine-grained semantic map as input $\mathcal{M}_{1}\in\mathbb{R}^{N\times H_{1}W_{1}}$ to the following Multi-level Mask Aggregation module.
Figure 3: Structure of STM and STQ module. STM takes visual and linguistic features as input and enables bi-directional modality calibration to generate the corresponding fine-grained semantic map at each level. Subsequently, STQ transfers modality complementary contents at the query level with the assistance of the referring semantics and original linguistic information. Further, we visualize the semantic map at each level, showcasing the semantic alignment at different granularities.
Semantic To Query Module.
The textual description in GRES weighs more crucial than the classic RES task in two folds: 1) It additionally encompasses implicit or complex multiple-target information such as counting expressions and relationships among various domains of instances. 2) The empty-target description appears relevant to the image, which will easily result in undesirable object responses. However, previous methods [20, 11] directly utilize the outputs from visual-attended cross attention, i.e., $\mathcal{F}^{L\sim V}$ for the consecutive cross-modal interaction. This approach potentially contributes to the loss of original linguistic information as contents in visual features are intensive but mostly irrelevant.
To address the issues, we propose the Semantic-to-Query (STQ) module, which performs comprehensive multi-modal complementary at the query level. It well incorporates the important referring semantics and preserves valuable language information simultaneously. Formally, we leverage fully connected layers to project the fine-grained semantic map $\mathcal{S}$ into query space:
@@
Then, as shown in Fig. 3, the high-level semantics in $\mathcal{Q}^{S}$, instead of the single visual modality, is injected into the language query $\mathcal{Q}^{L}$, i.e., the initial input query $\mathcal{L}^{O}$ or the one from previous STQ module:
@@
where $\mathrm{MHA}$ is the vanilla multi-head attention mechanism and $\mathcal{Q}^{L^{\prime}}\in\mathbb{R}^{N\times D}$.
As illustrated in Fig. 2, $\mathcal{Q}^{L^{\prime}}$ is further processed in the next-level STM module. It is anticipated that details within visual features become more adequate as the scale increases, which necessitates the more fine-grained modality interaction. Moreover, to prevent the shifted emphasis on original textual information and facilitate a deeper understanding of that with incorporated holistic semantics, we perform language reactivation by reintroducing the original linguistic feature $\mathcal{L}$. Thus, $\mathcal{Q}^{L^{\prime}}$ gets greatly enriched by the progressive modality complementary.
Multi-level Mask Aggregation.
Multiple-target referents typically vary in shape or scale within an image. The existing GRES mask decoding methods which obtain segmentation directly by leveraging the information-redundant joint feature, may lead to undesirable object response in multiple/non-target scenarios. Therefore, we propose Multi-level Mask Aggregation (MMA) to unify the semantic clue progressively for final mask predictions. In MMA, the corresponding object responses of semantic-enriched masks from different levels are hierarchically aggregated, which enhances the activation of proper regions while suppressing the unrelated ones.
Specifically, as shown in Fig. 2, the outputs $\{\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}\}$ from cascaded STMs are gathered into MMA for final segmentation. Firstly, we convert each level of semantic-enriched segmentation into a two-class mask by multiplying dynamic kernel $B$:
$\mathcal{M}_{i}^{\prime}=\mathcal{M}_{i}\times B_{i},$ (7)
where $B_{i}$ is conditional on $\mathcal{Q}^{L^{\prime}}_{i}$ from each STQ module. Subsequently, they are weighted aggregated through sequential upsampling in a multi-level paradigm, which can be expressed as:
@@
where $\mathrm{Up}(\cdot)$ indicates bilinear upsampling operation and $\alpha$ controls the fusion ratio between levels.
3.3 Multi-Adaptive Object Counting
Existing GRES methods tend to directly adopt a binary classification branch to judge the existence of objects. However, we observe that such simple design is not completely applicable in general referring segmentation settings. Particularly, it broadly categorizes multiple-target and single-target scenarios into one objective, ignoring their inherent specificity, which can easily incur ambiguity.
Meanwhile, several studies [21, 11] use the language reconstruction technique. Yet, it is not suitable in GRES since reconstructing unrelated content increases the burden of precise object awareness. Therefore, we propose Multi-Adaptive Object Counting (MAC) to endow our model with category-specific counting ability in multiple/single/non-target situations, which facilitates the understanding of both complicated sentences, e.g., counting, compound structure expression, and visual scenes.
Formally, the multi-modal semantic queries $\{\mathcal{Q}^{L^{\prime}}_{1},\mathcal{Q}^{L^{\prime}}_{2},\mathcal{Q}^{L^{\prime}}_{3}\}$ corresponding to each grain are transformed into the counting vector of $C$ categories, respectively:
$\mathcal{C}_{i}=\mathrm{MLP}(\mathcal{Q}^{L^{\prime}}_{i}),$ (9)
where $\mathcal{C}_{i}\in\mathbb{R}^{N\times C}$. Then $\{\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}$ are incorporated for the counting prediction $\mathcal{C}_{all}\in\mathbb{R}^{N\times C}$ by average pooling. Finally, we provide the explicit supervision on $\mathcal{C}_{all}$ by computing the $\mathit{L_{1}}$ distance between the multi-level accumulated results $\mathcal{C}^{pred}$ and the ground truth $\mathcal{C}^{gt}$, where $\mathcal{C}^{gt}$ (see in Fig. 2) is annotated as the exact counts of each category. The whole process can be formulated as:
[…]
On one hand, the MAC module is designed to seamlessly adapt to the specificity in generalized scenes, i.e., multiple/single/non-target tasks, which benefits the general object matching. On the other hand, guided by the the precise categories and counts of referent targets, our method facilitates comprehensive object perception in both image and language wise.
3.4 Output and Loss
[…]
Therefore, total loss with balance factor $\lambda$ can be formulated as:
[…]
4 Experiment
4.1 Experiment Setups
We evaluate our model on three GRES benchmarks using official evaluation metrics: gRefCOCO [20], R-RefCOCO/+/g [30] and Ref-ZOM [11]. Following previous GRES methods, we also conduct experiments on the RES dataset: RefCOCO/+/g [15, 27]. Elaborate descriptions of each benchmark and metrics are in Sec. A.1, Sec. A.2. We adopt Swin Transformer [25] and BERT [5] as the backbone. Due to space limitations, complete implementation details are in Sec. A.3.
Generalized Referring Expression Segmentation on gRefCOCO
@@ 4.1 Experiment Setups
HDC (Ours) Swin-B 68.28 65.42 63.38 72.52 71.60 65.29 63.85 62.79 60.68
Table 1: Comparison with the state-of-the-art methods on gRefCOCO. -V-7B means Vicuna-7B. (ft) denotes that the model is finetuned on the training set of gRefCOCO. denotes that the results are from the official repository.

### v2

# CoHD: A Counting-Aware Hierarchical Decoding Framework for Generalized Referring Expression Segmentation

@@ Abstract
The newly proposed Generalized Referring Expression Segmentation (GRES) amplifies the formulation of classic RES by involving complex multiple/non-target scenarios. Recent approaches address GRES by directly extending the well-adopted RES frameworks with object-existence identification. However, these approaches tend to encode multi-granularity object information into a single representation, which makes it difficult to precisely represent comprehensive objects of different granularity. Moreover, the simple binary object-existence identification across all referent scenarios fails to specify their inherent differences, incurring ambiguity in object understanding. To tackle the above issues, we propose a Counting-Aware Hierarchical Decoding framework (CoHD) for GRES. By decoupling the intricate referring semantics into different granularity with a visual-linguistic hierarchy, and dynamic aggregating it with intra- and inter-selection, CoHD boosts multi-granularity comprehension with the reciprocal benefit of the hierarchical nature. Furthermore, we incorporate the counting ability by embodying multiple/single/non-target scenarios into count- and category-level supervision, facilitating comprehensive object perception. Experimental results on gRefCOCO, Ref-ZOM, R-RefCOCO, and RefCOCO benchmarks demonstrate the effectiveness and rationality of CoHD which outperforms state-of-the-art GRES methods by a remarkable margin. Code is available at here.
00footnotetext: 🖂 Corresponding author.
@@ 1 Introduction
Figure 1: Comparison of decoding paradigms. Previous GRES methods attempt to integrate multi-granularity context into one joint feature for mask and object-existence prediction (a), which exhibits vulnerability to the complexity of multiple/non-target scenarios (c). Instead of relying on the single representation, we decouple the referring correspondence into the semantic hierarchy of different granularity and perform dynamic selective aggregation for robust segmentation. Further, in contrast to simple binary classification, we empower CoHD with counting ability to promote object perception (b).
@@
However, it poses new challenges with the additional generalized settings. The multiple-target referent necessitates more comprehensive semantic clues to handle complicated spatial-text relationships between instances for arbitrary location segmentation. On the other hand, the empty-target one highlights the significance of contextual information in mitigating undesirable responses from similar objects. An intuitive approach is to endow prevalent RES methods [36, 5, 6] with multi/non-target referring capability. For instance, ReLA [22] introduces a large-scale generalized dataset gRefCOCO and utilizes spatial-aware queries to transform GRES into a related region retrieval task. Meanwhile, DMMI [10] and RefSegformer [33] enhance modality interaction in the early-fusion process and incorporate new benchmarks Ref-ZOM and R-RefCOCO respectively, to realize robust segmentation.
Despite advancements in GRES, the direct extension from RES to GRES exhibits less compatibility in two aspects: 1) Incomplete Semantic Context: as shown in Fig. 1 (a) and (c), the previous decoding paradigm is solely dependent on a single representation for multi-granularity modeling, which leads to incomplete or incorrect object responses. We speculate that primarily stems from the increased complexity of referring semantics in GRES, e.g., intricate spatial relationships among instances, and deceptive expressions. It easily incurs the shifted emphasis on both visual and language-wise, thereby impeding the integration of all-grained contexts into one joint representation. 2) Imprecise Object Perception: previous GRES methods adopt simple binary classification for object-existence identification, which ignores the inherent distinction between multiple and single referent scenarios, potentially incurring ambiguity in object understanding.
To well address the challenges of GRES and break through the limitations of previous methods, we delve into a further investigation. Firstly, the extra referring scenarios, i.e., multiple/empty-target referents, bring more difficulty in accurately activating complete regions and discerning the non-related ones. For instance, as shown in Fig. 1 (b), different referring object activations vary in shapes, scales, and locations, and misinterpretation potentially occurs at the latter level. It indicates that simply relying on a single joint feature for mask decoding is insufficient. Instead, the complex referring semantics should be decoupled by different granularity and aggregated hierarchically for robust segmentation. Secondly, the inherent distinction of multiple/single referent cases should not be ignored since it makes a difference to the precision object perception when encountering implicit expressions such as “A line of cars”. Inspired by that, we present a Counting-aware Hierarchical Decoding framework (CoHD). Specifically, a multi-grained visual-linguistic hierarchy is formulated by the Hierarchical Semantic Decoder (HSD). Instead of simply leveraging a single joint feature, HSD performs dynamic hierarchical aggregation with intra- and inter-selection. Thus, the intricate referring semantic clues can be completely leveraged. Furthermore, we introduce the Adaptive Object Counting module (AOC) to specify the inherent distinction among multiple/single/non-target scenarios by embodying each into count- and category-level supervision, which tackles ambiguity and facilitates precise object perception.
We conduct extensive experiments on three popular GRES benchmarks, i.e., gRefCOCO [22], Ref-ZOM [10] and R-RefCOCO [33] along with RES datasets RefCOCO/+/g [15, 29]. Experimental results demonstrate that CoHD remarkably outperforms previous works on all benchmarks, showcasing the effectiveness of the proposed method for complicated scenarios in GRES.
@@
  • We present the Counting-Aware Hierarchical Decoding framework for GRES, aiming to decouple the complex referring semantics into different levels for boosting multi-granularity comprehension. In CoHD, the Hierarchical Semantic Decoder (HSD) formulates a visual-linguistic hierarchy of different granularity and performs mask decoding by dynamic hierarchical aggregation.
  • We introduce the Adaptive Object Counting module (AOC) to facilitate precise object perception by assigning explicit counting capability, which seamlessly adapts to the specificity of multiple/single/non-target scenarios.
  • Our approach outperforms all existing state-of-the-art methods by a remarkable margin without bells and whistles, e.g., $+4.8\%$ gIoU in gRefCOCO, $+7.5\%$ rIoU in R-RefCOCO and $+6.3\%$ Acc. in Ref-ZOM.
@@ 2 Related Work
Figure 2: Overview of CoHD. The model takes an image with the referring expression as input. After the encoding process, the Hierarchical Semantic Decoder (HSD) decouples the referring semantics into different granularity for establishing visual-linguistic hierarchy and then performs dynamic aggregation with intra- and inter-selection for final segmentation. Moreover, with well-aligned semantic contexts, we introduce the Adaptive Object Counting module (AOC) to fully promote comprehensive object perception with the category-specific counting ability under multiple/single/non-target scenarios.
@@
Referring Expression Segmentation (RES) aims to localize a single object referred by a natural language description, which is first introduced by Hu et al. [9]. Conventional approaches [9, 21, 1, 11, 6] typically utilize convolution-based operations as the cross-modal fusion to generate segmentation masks. To overcome the weakness of insufficiency in modeling visual-text relation that prior works suffer, recent studies [36, 5, 23, 26, 16] design advanced attention-based mechanism [31] to facilitate multi-modal interaction. Since previous works are restricted by pre-defined rules that one sentence must match an object, ReLA [22] proposes the Generalized Referring Expression Segmentation (GRES) task, which additionally involves empty-target and multiple-target scenarios. DMMI [10] introduces a new benchmark and baseline to realize beyond one target segmentation. Similarly, RefSegformer [33] endows the transformer-based model with empty-target sentence discrimination. Further, GSVA [35] reuses the [SEG] token and additionally incorporates the [REJ] one into LISA [18], which equips the MLLM with GRES ability. However, these methods seek to encode multi-granularity object information into a single joint representation for mask prediction, which is impractical in GRES. In contrast, we aggregate visual-linguistic semantic hierarchy of different granularity in a hierarchical manner for more generalized segmentation.
@@
Object counting is to predict the number of conditional objects within an image, which can be divided into detection-based and regression-based approaches. The detection-based methods [12, 7, 8] obtain the number of instances by proposals localization and ranking while the regression-based ones [2, 19, 30] generates predicted counts via density map integration. Different from the classic object counting mechanism, our proposed Adaptive Object Counting module performs at the query level. By considering various scenarios, i.e., multiple/single/non-target referents, our model achieves comprehensive object perception under each situation.
@@ 3.1 Overview
Figure 3: Illustration of Hierarchical Semantic Decoder. The Semantic Decoding Module takes visual and linguistic features as input and enables bi-directional modality calibration to generate the corresponding fine-grained semantic map at each level, which formulates a visual-linguist hierarchy. Subsequently, we incorporate the referring semantics into the query level with the original language information, which facilitates a more holistic object understanding. With the well-aligned semantic hierarchy, we perform dynamic aggregation with the intra- and inter-selection mechanism for mask decoding to excavate the potential of reciprocal benefits brought by the hierarchical nature.
@@
Fig. 2 depicts the overall pipeline of CoHD. Given an image $\mathcal{I}$, the visual backbone, e.g., Swin-Base [27] is to extract a set of vision features $\mathcal{V}_{i}\in\mathbb{R}^{D^{V}\times H_{i}\times W_{i}}$, $i\in\{1,2,3,4\}$, where $H_{i}$ and $W_{i}$ denote the width and height of each scale of feature maps respectively. $D^{V}$ signifies the channel dimension. $\mathcal{V}_{i}$ is further encoded as mask features $\mathcal{V}^{m}_{i}$ ($i\in\{1,2,3\}$) by deformable attention [38]. Simultaneously, the textual expression $\mathcal{E}$ with $T$ words is sent into a transformer-based [4] language encoder which generates the word-level embedding $\mathcal{L}^{w}\in\mathbb{R}^{T\times D^{L}}$ and sentence-level embedding $\mathcal{L}^{s}\in\mathbb{R}^{D^{L}}$. Considering that $\mathcal{L}^{w}$ contains fine-grained descriptions while $\mathcal{L}^{s}$ expresses the general information, language query $\mathcal{L}^{O}\in\mathbb{R}^{N\times D^{L}}$ is obtained via the concatenation of them. $\mathcal{L}^{O}$ and $\mathcal{V}^{m}_{i}$ are simultaneously sent into the Hierarchical Semantic Decoder (HSD), which generates corresponding object segmentation $\mathcal{M}\in\mathbb{R}^{H\times W}$ by dynamic hierarchical mask decoding on the semantic hierarchy of different granularity. Subsequently, Adaptive Object Counting (AOC) is designed to promote the precise object perception of different scenarios by counting ability.
3.2 Hierarchical Semantic Decoder
Previous GRES methods [22, 10, 33] adopt the traditional RES decoding paradigm which only incorporates the last modality-fused features. It is acceptable in RES since it is to match the single referred object through multi-modal interaction. However, it is less compatible with GRES. We observe that misinterpretation of region activation potentially occurs at the specific granularity in multiple-target cases, which indicates that simply performing sequential matching within each level as RES will lead to incomplete object segmentation. For the non-target referent, the final joint representation of multi-grained information is highly redundant, possibly resulting in incorrect object responses. Therefore, the hierarchical aggregation across granularity is significant for GRES, since it benefits from the reciprocal properties of semantic activation at each granularity, which helps to address the previous issues. Inspired by that, we propose a Hierarchical Semantic Decoder to formulate a visual-linguistic hierarchy through concatenated semantic decoding modules and perform hierarchical dynamic aggregation for robust segmentation.
Semantic Decoding Module.
The core idea of SDM is in two folds: 1) Mask Generation: generate the fine-grained semantic map at each scale of visual features which is further utilized in the dynamic aggregation process. 2) Query Refinement: the well-supervised semantic map helps guide the formulation of holistic comprehension of the referring semantics at the query level. Thus, the intricate semantics are well modeled by the mutual-enhanced relationship within modules.
@@ 3.2 Hierarchical Semantic Decoder
Specifically, taking $\mathcal{V}^{m}_{1}$ as an example, we firstly leverage fully connected layers to map visual features and language query into a shared joint space $D$, respectively. The mapped representations are further denoted as $\mathcal{V}$ and $\mathcal{L}$ for clarity. Then, we derive the coarse semantic map $\mathcal{A}$ by attention mechanism and it is applied in visual-linguistic interaction to align information flow between modalities for generating the fine-grained semantic map $\mathcal{S}$. The modality complementary highlights the corresponding object activation of language-guided visual features, while simultaneously relieving the potential noise in visual-oriented textual representations. The above process can be formulated as:
@@
$\mathcal{S}=\mathcal{F}^{L\sim V}\times(\mathcal{F}^{V\sim L})^{T},$ (4)
where $\mathcal{W}^{k}_{V}$, $\mathcal{W}^{v}_{V}$, $\mathcal{W}^{k}_{L}$, $\mathcal{W}^{v}_{L}$ are learnable weights that map visual and linguistic representations into attention space. In this manner, the key tokens related to the referent are anticipated to emerge significant responses in the semantic map $\mathcal{S}$ while suppressing the activation of non-related regions. Then, it is further utilized as the input $\mathcal{M}_{1}\in\mathbb{R}^{N\times H_{1}W_{1}}$ to formulate the multi-granularity visual-linguistic correspondence hierarchy for dynamic aggregation.
Considering that the textual description in GRES weighs more crucial than the classic RES task since it encompasses additionally implicit multiple-target information or deceptive description, we further introduce the well-aligned semantics into the query level for better object perception (see in Fig. 3). It well incorporates the important referring semantics and preserves valuable language information simultaneously. Formally, we leverage fully connected layers to project the fine-grained semantic map $\mathcal{S}$ into query space:
@@
The high-level semantics in $\mathcal{Q}^{S}$, instead of the single visual modality, is injected into the language query $\mathcal{Q}^{L}$, i.e., the initial input query $\mathcal{L}^{O}$ or the one from previous SD module:
@@
where $\mathrm{MHA}$ is the vanilla multi-head attention mechanism and $\mathcal{Q}^{L^{\prime}}\in\mathbb{R}^{N\times D}$. With greatly enriched by the referring semantics from $\mathcal{Q}^{S}$, we further perform language reactivation by reintroducing the original linguistic feature $\mathcal{L}$, which not only prevents the shifted emphasis but also facilitates a more holistic understanding of that.
Dynamic Hierarchical Aggregation.
We observe that the desired object responses typically vary in granularity when performing multiple object referring. Similar to the non-target one, the unrelated region activation will potentially occur due to the deceptive descriptions. Therefore, we emphasize the importance of reciprocal properties of multi-granularity information in GRES by proposing Dynamic Hierarchical Aggregation to unify the visual-linguistic hierarchy with intra- and inter-selective mechanisms. Shown in Fig. 3, the output $\{\left(\mathcal{M}_{1},\mathcal{Q}^{L^{\prime}}_{1}\right),\left(\mathcal{M}_{2},\mathcal{Q}_{2}^{L^{\prime}}\right),\left(\mathcal{M}_{3},\mathcal{Q}_{3}^{L^{\prime}}\right)\}$ from cascaded SDMs are gathered into DHA for segmentation. It is anticipated that $\mathcal{M}_{1}$ with the largest downsampling rate contains the highest level of object semantics, e.g., location, being referred or not, etc. In contrast, $\mathcal{M}_{3}$ mainly perceives the shape, and texture character of the objects. However, simply aggregation by summation easily introduces bias accumulation, leading to unsatisfactory outcomes. To fully excavate the benefit of hierarchical nature, we perform intra- and inter-selection on $\mathcal{M}$ and $\mathcal{Q}^{L^{\prime}}$ respectively. 1) Inter-Selection: inspired by the current success of MoE [13], we design a gating network to assign the saliency score for each level, which adaptively highlights the related object responses through granularity:
$\alpha_{i}=\mathcal{G}(\mathcal{Q}_{i}^{L^{\prime}}),i\in\{1,2,3\}$ (7)
$\mathcal{G}(x):=\mathrm{Sigmoid}(x\mathcal{W}_{g}).$ (8)
where $\alpha$in $\mathbb{R}^{1}$ and $\mathcal{W}_{g}$ is the learnable weights. 2) Intra-Selection: affected by the deceptive or compound structure textual description, some tokens may exhibit sparse and cluttered information pertaining to targets in $\mathcal{M}$. We assume that the informative and highly related region activation plays a significant role in hierarchical aggregation. With the purpose of dynamically enhancing the referred object responses and simultaneously suppressing the non-related ones, we leverage channel attention module [32] on $\mathcal{M}_{i}$ to evaluate the degree of referring correspondence in token level:
$\mathcal{M}_{i}^{\prime}=\mathrm{ChannelAttn}(\mathcal{M}_{i})\cdot\mathcal{M}_{i}.$ (9)
After suppressing the unrelated activation based on the low referring correspondence, $\{\mathcal{M}_{1}^{\prime},\mathcal{M}_{2}^{\prime},\mathcal{M}_{3}^{\prime}\}$ are weighted aggregated with each corresponding saliency score $\alpha_{i}$ through sequential upsampling in a multi-level paradigm. The whole process can be expressed as:
@@
Finally, we convert the aggregated semantic-enriched segmentation $\mathcal{M}^{*}_{3}$ into a two-class mask $\mathcal{M}^{o}$ by multiplying kernel $B$ which is conditional on the $\mathcal{Q}_{3}^{L^{\prime}}$:
$\mathcal{M}^{o}=\mathcal{M}^{*}_{3}\times B.$ (11)
@@
CoHD (Ours) Swin-B 68.42 65.17 63.68 72.67 71.85 64.00 63.60 62.63 60.37
Table 1: Comparison with the state-of-the-art methods on gRefCOCO. -V-7B means Vicuna-7B. (ft) denotes that the model is finetuned on the training set of gRefCOCO. denotes that the results are from the official repository. * indicates that experimental results using the same hyperparameters configuration from the official repository.